OpenStudy (scarlettfarra2000):

Will someone please help me question is below what you do is Add. Simplify by collecting like radical terms if possible

OpenStudy (scarlettfarra2000):

\[9\sqrt{27}+6\sqrt{48}\]

Mehek (mehek14):

simplify √27 and √48 first

OpenStudy (scarlettfarra2000):

I can do the first part \[9\sqrt{9}\sqrt{3}+6\sqrt{16}\sqrt{3}\] but i can't do the next part

Mehek (mehek14):

so 9√9 = 9*3 because √9 = 3 so 9*3 = 27 27√3 then 6√16 = 6*4 because √16 = 4 so 6*4 = 24 24√3 add 27√3 + 24√3 = ?

OpenStudy (scarlettfarra2000):

\[51\sqrt{3?}\]

Mehek (mehek14):

ye

OpenStudy (scarlettfarra2000):

Thank you so much @Mehek14

Mehek (mehek14):

yw :)